OpenStudy (anonymous):

Solve the following system of equations. 2x – 3y + z = –1 3x + 2y + 2z = –1 x – y – 3z = –4 (–1, 0, 1) (1, 1, 4) (–1, –1, –2) (1, 2, 1)

OpenStudy (anonymous):

Multiply the first by -2 and add to the second to get \[-4 x + 6 y - 2z = 2\\ 3x + 2y + 2z = –1 -x + 8y =1 \] Only the first (-1,0,1) satisfies -x+8 y=1

OpenStudy (anonymous):

\[ -4 x + 6 y - 2z = 2\\ 3x + 2y + 2z = –1\\ -x + 8y =1 \]

OpenStudy (anonymous):

You only can use this method if the question is that one of the choices is the solution. In this case, it is.